Compute the present value of a $3,600 deposit in year 2 and another $3,100 deposit at the end of year 4 if interest rates are 9 percent.
1
Expert's answer
2018-07-03T05:39:08-0400
The present value of a $3,600 deposit in year 2 is PV = 3,600/(1 + 0.09)^2 = $3,030.05
and another $3,100 deposit at the end of year 4 if interest rates are 9 percent is
PV = 3,100/(1 + 0.09)^4 = $2,196.12.